OpenStudy (swissgirl):

@safl;jk tan^2 x=sec^2x-1

OpenStudy (solomonzelman):

true :)

OpenStudy (p0sitr0n):

use sin^2x+cos^2x=1 , divide all by cos^2x

OpenStudy (swissgirl):

hahaha yaaaa We need to show @safl;jk how this is an identity What i would do is simplify the right hand side and show that it equals the left hand side

OpenStudy (swissgirl):

The following are 2 trig identities that will help you solve the question: \[ \sec^2x=\frac{1}{\cos^2x}\] \[ \sin^2x + \cos^2x =1\]

OpenStudy (swissgirl):

So lets only focus on the right hand side of the equation which is \( \sec^2x-1\) Right Hand Side: \[ \sec^2x -1\] \[ \frac{1}{\cos^2x}-1\] \[\frac{1}{\cos^2x} - \frac{ \cos ^2x }{\cos^2x}\] \[ \frac{1-\cos^2x}{\cos^2x}\] Now we know the trig identity that sin^2x+cos^2x =1 Using some algebra we can rearrange the equation to : 1-cos^2x =sin^2x Therefore we can replace 1-cos^2x with sin^2x in our equation above \[ \frac{1-\cos^2x}{\cos^2x}=\frac{\sin^2x}{\cos^2x} =\tan^x\] SO the right hand side of the equation is tan^2x and so the right hand side of the equation equals the left hand side of the equation and therefore we can agree that this equation is true

OpenStudy (swissgirl):

@safl;jk Come check this out